# Bushfires and Balance: Proactive versus Reactive Policies in Prescribed Burning

### v1

@@ 3.1. Some preliminary algebraic observations
$\triangle=(1+2c)^{2}c^{2}(2+c)^{2}\alpha^{6}-4c^{3}(2+c)^{3}\alpha^{6}+O(\alpha^{4})=c^{2}\Xi\alpha^{6},$
@@
$\Xi:=(1+2c)^{2}(2+c)^{2}-4c(2+c)^{3}+O\left(\frac{1}{\alpha^{2}}\right)=\Xi_{0}+O\left(\frac{1}{\alpha^{2}}\right).$
Hence, for $\alpha$ large enough (possibly depending on $\beta$, $\theta$, and $c$), we conclude that $\Xi\geqslant\frac{\Xi_{0}}{2}>0$ and accordingly $\triangle\geqslant\frac{c^{2}\Xi_{0}\alpha^{6}}{2}>0$.
@@
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)-\beta\eta|\theta|f_{0}=(1+2c)c(2+c)-c$
$\qquad=c\big{(}(1+2c)(2+c)-1\big{)}>c>0$
@@
$\triangle=-4(1+2c)^{3}c+O(c^{2})<0,$
as long as $c$ is sufficiently small.
@@
$\triangle=-27c^{2}\alpha^{2}\beta^{2}\theta^{2}+O(\alpha^{4})<0.$

### v2

@@ 4.1. Some preliminary algebraic observations
(A.10) $\triangle=(1+2c)^{2}c^{2}(2+c)^{2}\alpha^{6}-4c^{3}(2+c)^{3}\alpha^{6}+O(\alpha^{4})=c^{2}\Xi\alpha^{6},$
@@
(A.11) $\Xi:=(1+2c)^{2}(2+c)^{2}-4c(2+c)^{3}+O\left(\frac{1}{\alpha^{2}}\right)=\Xi_{0}+O\left(\frac{1}{\alpha^{2}}\right).$
Hence, for $\alpha$ large enough (possibly depending on $\beta$, $\theta$, and $c$), we conclude that $\Xi\geqslant\frac{\Xi_{0}}{2}>0$ and accordingly $\triangle\geqslant\frac{c^{2}\Xi_{0}\alpha^{6}}{2}>0$ (see Figure 6 for a specific visualization of this argument).
@@
(A.12) $\begin{split}&\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)-\beta\eta|\theta|f_{0}=(1+2c)c(2+c)-c\\ &\qquad=c\big{(}(1+2c)(2+c)-1\big{)}>c>0\end{split}$
@@
(A.13) $\triangle=-4(1+2c)^{3}c+O(c^{2})<0,$
as long as $c$ is sufficiently small (see Figure 7 for a specific visualization of this argument).
Figure 7. Plot of $\triangle$ when $f_{0}:=1$, $\gamma:=\eta:=c\alpha$, $\zeta:=(1+c)\alpha$, $\alpha:=1$, $\theta:=\beta:=1$. Notice that $\triangle>0$ for small values of $c>0$.
@@
(A.14) $\triangle=-27c^{2}\alpha^{2}\beta^{2}\theta^{2}+O(\alpha^{4})<0.$